OpenStudy (anonymous):

A set of double six dominoes has 28 dominoes. If you draw 1 domino, what is the probability that you will draw the double six?

OpenStudy (anonymous):

1/28

OpenStudy (anonymous):

Can you explain your answer please?

OpenStudy (anonymous):

PROB= desired outcome/ all posible outcomes

OpenStudy (anonymous):

So it is 6 / 28?

OpenStudy (anonymous):

how many doubl sixes r there ?

OpenStudy (anonymous):

28

OpenStudy (anonymous):

i dont know that game, so tell me not the total, but the specified domino.

OpenStudy (anonymous):

It isn't a game, it is a math problem. My math book is so crazy... They are looking for the probablity that you will draw a double six out of 28.

OpenStudy (anonymous):

then its 1/28

OpenStudy (anonymous):

if there are 2 , then pob = 2=28=1/14 etc., ect,

OpenStudy (anonymous):

There is an add on: What are the odds in favor of you drawing the double six?

OpenStudy (anonymous):

2/28=1/14

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

Another add on: What are the odds against you drawing the double six?

OpenStudy (anonymous):

that i dont know

OpenStudy (anonymous):

Oh. Well, thank you for all of your help! :)

OpenStudy (anonymous):

kk